soft drink dispenser is designed to dispense 250 mL of drink per cup. However, there have




been recent reports to the managements of a fast-food restaurant of both underfilled and




overfilled cups. Seeking to investigate this matter, they fill up 50 cups using this machine. If the mean amount of drink in the 50 cups is 248 mL with standard deviation of 12 mL, is there a




cause for concern for the management? Use a level of significance of 0.01

1
Expert's answer
2022-05-24T09:08:35-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=250"

"H_1:\mu\not=250"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=49" and the critical value for a two-tailed test is "t_c = 2.679952."

The rejection region for this two-tailed test is "R = \{t:|t|>2.679952\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{248-250}{12/\sqrt{50}}=-1.1785"

Since it is observed that "|t|=1.1785<2.679952=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=49" degrees of freedom, "t=-1.1785" is "p=0.244289," and since "p=0.244289>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 250, at the "\alpha = 0.01" significance level.